kind sir,
I have recently been introduced to linear and quadratic equations from the T.I.M.E guide for ntse. It has not been explained in brief so I would kindly ask you to send me a report over the subject. I am a student of class 8 and is therefore new to this with a desire to succeed. i would also want to ask you a question that how would we obtain two roots negative and positive during the calculation of x+y = 10.
thank you

Answer by lenny460(1073)

kind sir,
I have recently been introduced to linear and quadratic equations from the T.I.M.E guide for ntse. It has not been explained in brief so I would kindly ask you to send me a report over the subject. I am a student of class 8 and is therefore new to this with a desire to succeed. i would also want to ask you a question that how would we obtain two roots negative and positive during the calculation of x+y = 10.
thank you

Linear and quadratic equations are very easy once you grasp their concepts.

x + y = 10
-4 + 14 = 10




Lennox Obuong
Algebra 2 Student
Email: obuong3@aol.com